Circular No. 7594

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4643 SAGITTARII M. Della Valle, Arcetri Astrophysical Observatory, Florence; L. Da

Silva, Observatorio Nacional, Rio de Janeiro; E. Pompei, European Southern Observatory; and R. Williams, Space Telescope Science Institute, report: “Preliminary analysis of a spectrum (range 390–900 nm, resolution 48000) of V4643 Sgr (Nova Sgr 2001), obtained on Feb. 26.35 UT at La Silla with the 1.5-m telescope (+ Fiber-fed Extended Range Optical Spectrograph), shows a nova close to maximum light that has a heavily absorbed continuum. The spectrum is dominated by broad emission lines of the Balmer series, split into two asymmetric components and flanked by P-Cyg profiles. Also prominent are emission lines due to Na I (mult. 1), He I (mult. 11), N II (mult. 3, 21), He II (1), and O I (4). The lack of Fe II lines and the large expansion velocities (the FWZI measured for Hα yields ∼ 10000 km/s) suggest that this nova belongs to the He/N spectroscopic class (see Williams 1992, A.J. 104, 725).”

COMET C/2000 W1 (UTSUNOMIYA-JONES) Further to the report on IAUC 7586, A. C. Gilmore reports that 3-

min unfiltered CCD images taken on Mar. 3.61 UT with the University of Canterbury’s Mount John Observatory 1-m f/7.7 telescope showed only a diffuse parabolic glow at the comet’s expected position. The glow was brighter and ∼ 1′ across at the ‘head’ end. The ‘tail’, in p.a. 80^{o}, was at least 10′ long and widened to ∼ 2′ across at the frame’s edge. No stellar central condensation was found, though anything brighter than red mag 20 should have been detected.

COMET C/1999 T1 (McNAUGHT-HARTLEY) C. E. Woodward, J. E. Lyke, and R. D. Gehrz, University of Minnesota

(UM), report 7- to 23-μm photometry of this comet on Feb. 21.51 UT at the Mt. Lemmon Observing Facility 1.52-m telescope (+ UM bolometer + IRTF narrowband ‘silicate’ filters). No evidence for strong silicate emission (cf. IAUC 7582) was observed at 11 μm; a blackbody fit to the observed spectral energy distribution yields a color temperature of 270 ± 20 K. Observed magnitudes: [7.9 μm] = 3.19 ± 0.25, [8.8 μm] = 3.24 ± 0.22, [9.8 μm] = 2.93 ± 0.27, [10.3 μm] = 2.91 ± 0.12, [11.7 μm] = 1.68 ± 0.12, [12.5 μm] = 2.05 ± 0.22, [18.3 μm] = 0.42 ± 0.23, and [23.0 μm] = 0.62 ± 0.30.

2001 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green